One simple model for a person running the 100 mdash is to assume the sprinter runs with constant acceleration until reaching top speed, then maintains that speed through the finish line.

If a sprinter reaches his top speed of 11.5 m/sin 2.74 s , what will be his total time?
Answer #1

10 t assumeんese sprinter puts he moves wth Distanee ひし, time 5 ID0 눈,(2:74)(11.5)十@5 x 4 t) 100 = DD 5755 l5 So tálad time
Answer #2

SOLUTION :


Initial speed = 0 m/s

Top speed  = 11.5 m/s 


As top speed is attained at constant acceleration, we can consider average speed 

during the acceleration period to determine the distance run.


So, 


Distance run in the acceleration period

= average speed * time taken

= (0 + 11.5 / 2 * 2.74

= 5.75 * 2.74

= 15.755  m 


So, distance run at top speed of 11.5 m/s 

= 100 - 15.755

= 84. 245 m 


Time taken to `run at constant top speed

= Distance run / top sped

= 84.245 / 11.5

= 7.33 sec


Total time taken to run 100m race

= time at constant acceleration + time at constant sped

= 2.74 + 7.33

= 10.07 sec (ANSWER).

answered by: Tulsiram Garg